# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 3 Coderivative-Based Newton Method for Subgradient Systems
where $(\widehat{x}^{k},\widehat{v}^{k})\in\mbox{\rm gph}\,\partial\varphi$. This general framework plays a crucial role in designing and justifying our new coderivative-based Newton-type algorithms for solving nonsmooth and nonconvex optimization problems in Sections 4 and 6. Based on (3.3), a key aspect in the development and validation of our Newtonian algorithms is to establish the existence of and explicitly determine a direction $d\in{\rm I\!R}^{n}$ that satisfies the generalized Newton system (or the second-order subdifferential inclusion)
$-v\in\partial^{2}\varphi(x,v)(d)$ (3.4)
for a given $(x,v)\in\mbox{\rm gph}\,\partial\varphi$. Efficient conditions for solvability of (3.4) presented in [24, Theorems 3.1 and 3.2] ensure that the proposed algorithm is well-defined. Observe that the local closedness assumption for the limiting subdifferential graph imposed below holds in very general settings, especially when $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is subdifferentially continuous; see Section 2. The following results are drawn from [24, Theorems 3.1 and 3.2].
@@
However, Proposition 3.1 only provides a condition that guarantees the existence of a direction $d\in{\rm I\!R}^{n}$ satisfying (3.4) without offering explicit methods to find this direction. In practice, an explicit calculation of such a direction is crucial for the implementation of our algorithms. Let us now obtain new sufficient conditions that not only guarantee the existence of the required direction but also furnish a constructive procedure to calculate it for some important classes of functions. Before establishing the theorem, we present a preliminary result from [14, Proposition 4.2] needed in the proof.
@@
$\text{\rm gph}\,\partial q\subset\text{\rm gph}\,\partial^{2}\varphi(\bar{x},\bar{v})=\text{\rm gph}\,D^{*}\partial\varphi(\bar{x},\bar{v}).$ (3.5)
@@
$({\rm I\!R}^{n}\times\{0\})\cap\text{\rm gph}\,\partial q\subset({\rm I\!R}^{n}\times\{0\})\cap\text{\rm gph}\,D^{*}\partial\varphi(\bar{x},\bar{v})=\{(0,0)\}.$
Note that $\text{\rm gph}\,\partial q$ is a linear subspace of ${\rm I\!R}^{2n}$ since $q$ is a generalized quadratic form. By Lemma 3.2, we find a unique $n\times n$ matrix $C$ such that $Ce_{i}\in\partial q(e_{i})$ for all $i=1,\ldots,n$. This implies that $\partial q(w)\neq\emptyset$ whenever $w\in{\rm I\!R}^{n}$ since $\text{\rm gph}\,\partial q$ is a linear subspace of ${\rm I\!R}^{2n}$. Pick further any $u_{i}\in\partial q(e_{i})$ for $i=1,\ldots,n$ and define the $n\times n$ matrix $A(\bar{x},\bar{v})$ such that its $i$th column is $u_{i}$. To verify (i), observe that $w\in{\rm I\!R}^{n}$ satisfies the system of linear equations $A(\bar{x},\bar{v})w=0$ if and only if $\sum_{i=1}^{n}w_{i}u_{i}=0$. It follows from (3.5) that
@@
[Proof.] It follows from Proposition 2.3 that $\partial\varphi$ is locally strongly monotone around $(\bar{x},\bar{v})$ as in (2.15). Using [48, Corollary 6.3] and [55, Proposition 13.5], we deduce that $d^{2}\varphi(\bar{x},\bar{v})$ is an l.s.c. strongly convex function, and thus (3.6) admits a unique optimal solution $d\in{\rm I\!R}^{n}$. By [40, Proposition 1.30], we have the inclusion
@@
Employing the fundamental result of [55, Theorem 13.40] gives us the relationship between second subderivatives and graphical derivatives as follows:
@@
Moreover, it follows from [55, Theorem 13.57] that
@@
In general, solving the second-order subdifferential inclusion (3.4) may be a challenging issue. Nevertheless, our approach offers computational advantages in some rather broad specific scenarios. Consider the following points:
@@
(ii) Theorem 3.4 enables us to obtain a solution to (3.4) by solving a strongly convex optimization subproblem derived from second subderivatives. Calculus rules for second subderivatives have been thoroughly explored in [36, 55] for practical and broader classes of functions overwhelmingly arising in optimization and variational analysis. Moreover, when the cost function is generalized twice differentiable, the subproblem in (3.6) reduces to a quadratic program over a subspace, for which many efficient numerical algorithms are widely available.
(iii) When the cost function is the sum of two nonconvex functions, which is the primary focus of this paper, Theorem 4.11 offers an alternative approach to finding the Newton direction by computing the Bouligand Jacobian of the proximal mapping $\partial_{B}\text{\rm Prox}_{\lambda g}$. Explicit formulas for these computations are provided in, e.g., [46, Section 5] covering a wide range of important optimization problems.
(iv) In some situations, advantages of the second-order subdifferential inclusion (3.4) in coderivative-based Newtonian methods become particularly clear when compared to solving linear equations using the Clarke Jacobian in semismooth Newton methods. This distinction is highlighted in [27, Section 6.1]. In such cases, the direction generated through (3.4) has an explicit formula, while semismooth Newton methods require solving a linear system. Computational steps in coderivative-based Newton algorithms depend solely on vector computations without the need for matrix calculus. This approach offers significant benefits in terms of computational cost. Our future goal is to identify practically important optimization problems, where the generalized Newton directions generated by (3.4) have explicit formulas. We aim to focus on cases where all computational costs involve only vector computations, leading us to a significant reduction in the computational costs of algorithms.
@@
$-\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$ (3.11)
@@
It follows from [39, Corollary 2.29] that the collection of points $(x,v)\in\mbox{\rm gph}\,\partial\varphi$ with $\partial\varphi(x)\neq\emptyset$ is dense in the subdifferential graph, and thus we can always select $(\widehat{x}^{k},\widehat{v}^{k})\in\mbox{\rm gph}\,\partial\varphi$ arbitrarily close to $(\bar{x},0)$. The issue is about the fulfillment of estimate (3.10). Let us show how to do it in the following major settings without knowing the solution $\bar{x}$.
@@
(ii) If $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is continuously prox-regular at $\bar{x}$ for 00 with modulus $r>0$, it follows from Proposition 2.5 that for each $\lambda\in(0,r^{-1})$, there is a neighborhood $U_{\lambda}$ of $\bar{x}$ such that $\text{\rm Prox}_{\lambda\varphi}$ is single-valued and Lipschitz continuous with modulus $L_{\lambda}>0$ on $U_{\lambda}$. Having $x^{k}\in U_{\lambda}$ at the $k^{\text{th}}$ iteration, choose $\widehat{x}^{k}$ and $\widehat{v}^{k}$ as
$\widehat{x}^{k}:=\text{\rm Prox}_{\lambda\varphi}(x^{k})\quad\text{and }\;\widehat{v}^{k}:=\frac{1}{\lambda}\left(x^{k}-\text{\rm Prox}_{\lambda\varphi}(x^{k})\right).$ (3.12)
@@
Recall from [43, Proposition 2.4] that for $F:=\nabla\varphi$, where $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, the semismoothness of $F$ at $(\bar{x},\nabla\varphi(\bar{x}))$ is equivalent to
@@
Let $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ be of class $\mathcal{C}^{1,1}$ around $\bar{x}\in{\rm I\!R}^{n}$. If $\nabla\varphi$ is metrically regular around $(\bar{x},\nabla\varphi(\bar{x}))$, then there exist $c>0$ and a neighborhood $U$ of $\bar{x}$ such that for any $x\in U$ and $d\in{\rm I\!R}^{n}$ satisfying $-\nabla\varphi(x)\in\partial^{2}\varphi(x)(d)$, we find $v_{x}\in\partial^{2}\varphi(x)(\bar{x}-x)$ with
@@
Hence for any $x\in U$ and $d\in{\rm I\!R}^{n}$ satisfying $-\nabla\varphi(x)\in\partial^{2}\varphi(x)(d)$, we find $v_{x}\in\partial^{2}\varphi(x)(\bar{x}-x)$ with
$-\nabla\varphi(x)-v_{x}\in\partial^{2}\varphi(x)(x+d-\bar{x}).$
@@
In the setting of Proposition 2.5, take any $\lambda\in(0,1/r)$, $x\in U_{\lambda}$ with $v=\nabla e_{\lambda}\varphi(x)$. Then we have the equivalence
@@
Let $\bar{x}$ solve (3.1), where $\partial\varphi$ is metrically regular around $(\bar{x},0)$ and semismooth at this point. Assume further that $\varphi$ is continuously prox-regular at $\bar{x}$ for 00. Then for any $\varepsilon>0$, there exists a neighborhood $W$ of $(\bar{x},0)$ such that
$\|x+d-\bar{x}\|\leq\varepsilon\|(x,v)-(\bar{x},0)\|$ (3.18)
whenever $(x,v)\in\text{\rm gph}\partial\varphi\cap W$ and $d\in{\rm I\!R}^{n}$ satisfy $-v\in\partial^{2}\varphi(x,v)(d)$.
[Proof.] Proposition 2.5 tells us that the assumed prox-boundedness and continuous prox-regularity properties of $\varphi$ guarantee the existence of $\lambda\in(0,1)$ such that there is a neighborhood $U_{\lambda}$ of $\bar{x}$ on which $e_{\lambda}\varphi$ is of class $\mathcal{C}^{1,1}$, that the proximal mapping $\text{\rm Prox}_{\lambda\varphi}$ is single-valued and Lipschitz continuous on $U_{\lambda}$ with ${\rm Prox}_{\lambda\varphi}(\bar{x})=\bar{x}$, and that the gradient expression for the Moreau envelope (2.19) holds. Since $\partial\varphi$ is semismooth at $(\bar{x},0)$, it follows from [43, Theorem 6.2, Claim 2] that $\nabla e_{\lambda}\varphi$ is semismooth at $(\bar{x},0)$. Given $\varepsilon>0$, we find a neighborhood $U$ of $\bar{x}$ on which we get by Lemma 3.8 that
$\|x^{\prime}+d^{\prime}-\bar{x}\|\leq\frac{\varepsilon}{2}\|x^{\prime}-\bar{x}\|\;\text{ if }\;x^{\prime}\in U\;\mbox{ and }\;d^{\prime}\in{\rm I\!R}^{n}\;\text{ satisfy }\;-\nabla e_{\lambda}\varphi(x^{\prime})\in\partial^{2}e_{\lambda}\varphi(x^{\prime})(d^{\prime}).$ (3.19)
@@
$-v\in\partial^{2}\varphi(x,v)(d)=\partial^{2}\varphi((x+\lambda v)-\lambda v,v)(d).$ (3.21)
@@
$-\nabla e_{\lambda}\varphi(x+\lambda v)=-v\in\partial^{2}e_{\lambda}\varphi(x+\lambda v,v)(d-\lambda v).$
This implies in turn due to (3.19) and (3.20) that
$\|(x+\lambda v)+(d-\lambda v)-\bar{x}\|\leq\frac{\varepsilon}{2}\|x+\lambda v-\bar{x}\|.$
@@
$\|x+d-\bar{x}\|\leq\frac{\varepsilon}{2}\left(\|x-\bar{x}\|+\lambda\|v\|\right)\leq\varepsilon\|(x,v)-(\bar{x},0)\|,$
which completes the proof of theorem.
@@
Let $\bar{x}$ solve (3.1), where $\partial\varphi$ is metrically regular around $(\bar{x},0)$ and semismooth at this point. Suppose that $\varphi$ is continuously prox-regular at $\bar{x}$ for 00. Then there is a neighborhood $U$ of $\bar{x}$ such that, whenever $x^{0}\in U$, Algorithm 1 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ that $Q$-superlinearly converges to $\bar{x}$ as $k\to\infty$.
[Proof.] Due to the robustness of the continuous prox-regularity of $\varphi$ at $\bar{x}$ for 00 and the metric regularity of $\partial\varphi$ around $(\bar{x},0)$, Theorem 3.3 implies that for each $(x,v)\in\mbox{\rm gph}\,\partial\varphi$ near $(\bar{x},\bar{v})$, there exists a direction $d\in{\rm I\!R}^{n}$ satisfying the generalized Newton system (3.4). Combining this with Theorem 3.10 allows us to find such $\delta>0$ that for any $x\in\mathbb{B}_{\delta}(\bar{x})$ there exists $(\widehat{x},\widehat{v})\in\mbox{\rm gph}\,\partial\varphi$ satisfying the estimate
@@
$-v\in\partial^{2}\varphi(x,v)(d)\;\text{and }\|x+d-\bar{x}\|\leq\frac{1}{2\eta}\|(x,v)-(\bar{x},0)\|.$ (3.23)
@@
which readily implies that $\left(\widehat{x}^{0},\widehat{v}^{0}\right)\in\mbox{\rm gph}\,\partial\varphi\cap\mathbb{B}_{\delta}(\bar{x},0)$. Furthermore, by (3.23) we find $d^{0}\in{\rm I\!R}^{n}$ satisfying $-\widehat{v}^{0}\in\partial^{2}\varphi(\widehat{x}^{0},\widehat{v}^{0})(d^{0})$, and hence
@@
This clearly implies that both sequences $\{x^{k}\}$ and $\{\widehat{x}^{k}\}$ converge to the solution $\bar{x}$ of (3.1) while the sequence $\{\widehat{v}^{k}\}$ converges to 00 as $k\to\infty$. To verify finally the $Q$-superlinear convergence of the iterative sequence $\{x^{k}\}$, pick any $\varepsilon>0$ and deduce from Theorem 3.10 that

### v3

@@ 3 Coderivative-Based Newton Method for Subgradient Systems
where $(\widehat{x}^{k},\widehat{v}^{k})\in\mbox{\rm gph}\,\partial\varphi$. This general framework plays a crucial role in designing and justifying our new coderivative-based Newton-type algorithms for solving nonsmooth and nonconvex optimization problems in Sections 4 and 6.
Based on (3.3), a key aspect in the development and validation of our Newtonian algorithms is to establish the existence of and explicitly determine a direction $d\in{\rm I\!R}^{n}$ that satisfies the generalized Newton system (or the second-order subdifferential inclusion)
$v\in\partial^{2}\varphi(x,v)(d)$ (3.4)
for a given $(x,v)\in\mbox{\rm gph}\,\partial\varphi$. Efficient conditions for solvability of (3.4) presented in [24, Theorems 3.1 and 3.2] ensure that the proposed algorithm is well-defined. Observe that the local closedness assumption for the limiting subdifferential graph imposed below holds in very general settings, especially when $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is subdifferentially continuous; see Section 2. The following results are taken from [24, Theorems 3.1 and 3.2].
@@
Proposition 3.1 only provides a condition that guarantees the existence of a direction $d\in{\rm I\!R}^{n}$ satisfying (3.4) without offering explicit methods to find this direction. In practice, an explicit calculation of such a direction is crucial for the implementation of our algorithms. Let us now obtain new sufficient conditions that not only guarantee the existence of the required direction but also furnish a constructive procedure to calculate it for some important classes of functions. Before establishing the theorem, we present a preliminary result from [14, Proposition 4.2] needed in the proof.
@@
$\mbox{\rm gph}\,\partial q\subset\mbox{\rm gph}\,\partial^{2}\varphi(\bar{x},\bar{v})=\mbox{\rm gph}\,D^{*}\partial\varphi(\bar{x},\bar{v}).$ (3.5)
@@
$({\rm I\!R}^{n}\times\{0\})\cap\mbox{\rm gph}\,\partial q\subset({\rm I\!R}^{n}\times\{0\})\cap\mbox{\rm gph}\,D^{*}\partial\varphi(\bar{x},\bar{v})=\{(0,0)\}.$
Note that $\mbox{\rm gph}\,\partial q$ is a linear subspace of ${\rm I\!R}^{2n}$ since $q$ is a generalized quadratic form. By Lemma 3.2, we find a unique $n\times n$ matrix $C$ such that $Ce_{i}\in\partial q(e_{i})$ for all $i=1,\ldots,n$. This implies that $\partial q(w)\neq\emptyset$ whenever $w\in{\rm I\!R}^{n}$ since $\mbox{\rm gph}\,\partial q$ is a linear subspace of ${\rm I\!R}^{2n}$. Pick further any $u_{i}\in\partial q(e_{i})$ for $i=1,\ldots,n$ and define the $n\times n$ matrix $A(\bar{x},\bar{v})$ such that its $i$th column is $u_{i}$. To verify (i), observe that $w\in{\rm I\!R}^{n}$ satisfies the system of linear equations $A(\bar{x},\bar{v})w=0$ if and only if $\sum_{i=1}^{n}w_{i}u_{i}=0$. It follows from (3.5) that
@@
[Proof.] It follows from Proposition 2.3 that $\partial\varphi$ is locally strongly monotone around $(\bar{x},\bar{v})$ in the sense that there are $s>0$ and a neighborhood $W$ of $(\bar{x},\bar{v})$ such that
$\langle v_{1}-v_{2},u_{1}-u_{2}\rangle\geq s\|u_{1}-u_{2}\|^{2}\quad\text{for all }\;(u_{1},v_{1}),(u_{2},v_{2})\in\mbox{\rm gph}\,\partial\varphi\cap W.$
Using [47, Corollary 6.3] and [54, Proposition 13.5], we deduce that $d^{2}\varphi(\bar{x},\bar{v})$ is an l.s.c. strongly convex function, and thus (3.6) admits a unique optimal solution $d\in{\rm I\!R}^{n}$. By [39, Proposition 1.30], we have the inclusion
@@
Employing the fundamental result of [54, Theorem 13.40] gives us the relationship between second subderivatives and graphical derivatives as follows:
@@
Moreover, it follows from [54, Theorem 13.57] that
@@
In general, solving the second-order subdifferential inclusion (3.4) may be a challenging issue. Nevertheless, our approach offers computational advantages in some rather broad scenarios.
@@
(ii) Theorem 3.4 enables us to obtain a solution to (3.4) by solving a strongly convex optimization subproblem derived from second subderivatives. Calculus rules for second subderivatives have been thoroughly explored in [35, 54] for practical and broad classes of functions overwhelmingly arising in optimization and variational analysis. Moreover, when the cost function is generalized twice differentiable, the subproblem in (3.6) reduces to a quadratic program over a subspace, for which many efficient numerical algorithms are widely available.
(iii) When the cost function is the sum of two nonconvex functions, which is the primary focus of this paper, Theorem 4.11 offers an alternative approach to finding the Newton direction by computing the Bouligand Jacobian of the proximal mapping $\partial_{B}\text{\rm Prox}_{\lambda g}$. Explicit formulas for these computations are provided in, e.g., [45, Section 5] covering a wide range of important optimization problems.
@@
$\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$
@@
It follows from [38, Corollary 2.29] that the collection of points $(x,v)\in\mbox{\rm gph}\,\partial\varphi$ with $\partial\varphi(x)\neq\emptyset$ is dense in the subdifferential graph, and thus we can always select $(\widehat{x}^{k},\widehat{v}^{k})\in\mbox{\rm gph}\,\partial\varphi$ arbitrarily close to $(\bar{x},0)$. The issue is about the fulfillment of estimate (3.10). Let us show how to do it in the following major settings without knowing the solution $\bar{x}$.
@@
(ii) If $\varphi\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is continuously $r$-level prox-regular at $\bar{x}$ for 0, it follows from Proposition 2.5 that for sufficiently small $\lambda>0$, there is a neighborhood $U_{\lambda}$ of $\bar{x}$ such that $\text{\rm Prox}_{\lambda\varphi}$ is single-valued and Lipschitz continuous with modulus $L_{\lambda}>0$ on $U_{\lambda}$. Having $x^{k}\in U_{\lambda}$ at the $k^{\text{th}}$ iteration, choose $\widehat{x}^{k}$ and $\widehat{v}^{k}$ as
[…]
@@
Recall from [42, Proposition 2.4] that for $F:=\nabla\varphi$, where $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, the semismoothness of $F$ at $(\bar{x},\nabla\varphi(\bar{x}))$ is equivalent to
@@
Let $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ be of class $\mathcal{C}^{1,1}$ around $\bar{x}\in{\rm I\!R}^{n}$. If $\nabla\varphi$ is metrically regular around $(\bar{x},\nabla\varphi(\bar{x}))$, then there exist $c>0$ and a neighborhood $U$ of $\bar{x}$ such that for any $x\in U$ and $d\in{\rm I\!R}^{n}$ satisfying $\nabla\varphi(x)\in\partial^{2}\varphi(x)(d)$, we find $v_{x}\in\partial^{2}\varphi(x)(\bar{x}-x)$ with
@@
Hence for any $x\in U$ and $d\in{\rm I\!R}^{n}$ satisfying $\nabla\varphi(x)\in\partial^{2}\varphi(x)(d)$, we find $v_{x}\in\partial^{2}\varphi(x)(\bar{x}-x)$ with
$\nabla\varphi(x)-v_{x}\in\partial^{2}\varphi(x)(x+d-\bar{x}).$
@@
In the setting of Proposition 2.5, take any $\lambda\in(0,\bar{\lambda})$, $x\in U_{\lambda}$ with $v=\nabla e_{\lambda}\varphi(x)$. Then we have the equivalence
@@
Let $\bar{x}$ solve (3.1), where $\partial\varphi$ is metrically regular around $(\bar{x},0)$ and semismooth at this point. Assume further that $\varphi$ is continuously prox-regular at $\bar{x}$ for 0. Then for any $\varepsilon>0$, there exists a neighborhood $W$ of $(\bar{x},0)$ such that
$\|x+d-\bar{x}\|\leq\varepsilon\|(x,v)-(\bar{x},0)\|$
whenever $(x,v)\in\mbox{\rm gph}\,\partial\varphi\cap W$ and $d\in{\rm I\!R}^{n}$ satisfy $v\in\partial^{2}\varphi(x,v)(d)$.
[Proof.] Proposition 2.5 tells us that the assumed prox-boundedness and continuous prox-regularity properties of $\varphi$ guarantee the existence of a sufficiently small $\lambda\in(0,1)$ such that there is a neighborhood $U_{\lambda}$ of $\bar{x}$ on which $e_{\lambda}\varphi$ is of class $\mathcal{C}^{1,1}$, that the proximal mapping $\text{\rm Prox}_{\lambda\varphi}$ is single-valued and Lipschitz continuous on $U_{\lambda}$ with ${\rm Prox}_{\lambda\varphi}(\bar{x})=\bar{x}$, and that the gradient expression for the Moreau envelope (2.15) holds. Since $\partial\varphi$ is semismooth at $(\bar{x},0)$, it follows from [42, Theorem 6.2, Claim 2] that $\nabla e_{\lambda}\varphi$ is semismooth at $(\bar{x},0)$. Given $\varepsilon>0$, we find a neighborhood $U$ of $\bar{x}$ on which we get by Lemma 3.8 that
$\|x^{\prime}+d^{\prime}-\bar{x}\|\leq\frac{\varepsilon}{\sqrt{2}}\|x^{\prime}-\bar{x}\|\;\text{ if }\;x^{\prime}\in U\;\mbox{ and }\;d^{\prime}\in{\rm I\!R}^{n}\;\text{ satisfy }\;-\nabla e_{\lambda}\varphi(x^{\prime})\in\partial^{2}e_{\lambda}\varphi(x^{\prime})(d^{\prime}).$ (3.17)
@@
$v\in\partial^{2}\varphi(x,v)(d)=\partial^{2}\varphi((x+\lambda v)-\lambda v,v)(d).$ (3.19)
@@
$\nabla e_{\lambda}\varphi(x+\lambda v)=-v\in\partial^{2}e_{\lambda}\varphi(x+\lambda v,v)(d-\lambda v).$
This ensures, with taking (3.17) and (3.18) into account, that
$\|x+d-\bar{x}\|=\|(x+\lambda v)+(d-\lambda v)-\bar{x}\|\leq\frac{\varepsilon}{\sqrt{2}}\|x+\lambda v-\bar{x}\|.$
@@
$\|(x,v)-(\bar{x},0)\|^{2}=\|x-\bar{x}\|^{2}+\|v\|^{2}\geq\|x-\bar{x}\|^{2}+\|\lambda v\|^{2}\geq\frac{1}{2}(\|x-\bar{x}\|+\|\lambda v\|)^{2},$
which imply in turn that
$\|x+d-\bar{x}\|\leq\frac{\varepsilon}{\sqrt{2}}\|x+\lambda v-\bar{x}\|\leq\frac{\varepsilon}{\sqrt{2}}\left(\|x-\bar{x}\|+\|\lambda v\|\right)\leq\varepsilon\|(x,v)-(\bar{x},0)\|$
and thus complete the proof of theorem.
@@
Let $\bar{x}$ solve (3.1), where $\partial\varphi$ is metrically regular around $(\bar{x},0)$ and semismooth at this point. Suppose that $\varphi$ is continuously prox-regular at $\bar{x}$ for 0. Then there is a neighborhood $U$ of $\bar{x}$ such that, whenever $x^{0}\in U$, Algorithm 1 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ that $Q$-superlinearly converges to $\bar{x}$ as $k\to\infty$.
[Proof.] Due to the robustness of the continuous prox-regularity of $\varphi$ at $\bar{x}$ for 0 and the metric regularity of $\partial\varphi$ around $(\bar{x},0)$, Theorem 3.3 implies that for each $(x,v)\in\mbox{\rm gph}\,\partial\varphi$ near $(\bar{x},\bar{v})$, there exists a direction $d\in{\rm I\!R}^{n}$ satisfying the generalized Newton system (3.4). Combining this with Theorem 3.10 allows us to find such $\delta>0$ that for any $x\in\mathbb{B}_{\delta}(\bar{x})$ there exists $(\widehat{x},\widehat{v})\in\mbox{\rm gph}\,\partial\varphi$ satisfying the estimate
@@
$v\in\partial^{2}\varphi(x,v)(d)\;\text{and }\|x+d-\bar{x}\|\leq\frac{1}{2\eta}\|(x,v)-(\bar{x},0)\|.$ (3.21)
@@
which readily implies that $\left(\widehat{x}^{0},\widehat{v}^{0}\right)\in\mbox{\rm gph}\,\partial\varphi\cap\mathbb{B}_{\delta}(\bar{x},0)$. Furthermore, by (3.21) we find $d^{0}\in{\rm I\!R}^{n}$ satisfying $\widehat{v}^{0}\in\partial^{2}\varphi(\widehat{x}^{0},\widehat{v}^{0})(d^{0})$, and hence
@@
This clearly implies that both sequences $\{x^{k}\}$ and $\{\widehat{x}^{k}\}$ converge to the solution $\bar{x}$ of (3.1) while the sequence $\{\widehat{v}^{k}\}$ converges to 0 as $k\to\infty$. To verify finally the $Q$-superlinear convergence of the iterative sequence $\{x^{k}\}$, pick any $\varepsilon>0$ and deduce from Theorem 3.10 that